OpenStudy (anonymous):

A kite 100 ft above the ground is moving horizontally at a speed of 8ft/s. At what rate is the angle between the string and the horizontal decreasing when 200 ft of string has been let out?

OpenStudy (nathan917):

the height of the kite h = 100 ft the horizontal distance of the kite = x = vt where v = 9 ft/s, t is time h, x and the string (length L) for a right triangle, the angle between the string and the horizontal A satisfies tan A = h/x sec²A dA/dt = -h/x² dx/dt we know dx/dt = v so dA/dt = -hvcos²A/x² = -hv/(x/cosA)² = -hv/L² when L = 200 ft, dA/dt = -100*9/200² = -0.0225 = -9/400 rad/s negative sign means decreasing.